# Polar Sparsity: High Throughput Batched LLM Inferencing with Scalable Contextual Sparsity

### v2

@@ C.1 Router Ablation Study
Figure 10 shows the decode latency of MLP and Attention blocks along with their respective routers, and includes the measured inference latency of sparse kernels across selected sparsity levels. The MLP router introduces approximately four times higher latency compared to the attention router. At higher activation levels, the combined latency of the MLP router and sparse MLP approaches, and in some cases exceeds, that of the dense MLP. To address this, we overlap the MLP router execution with the attention block, following a strategy similar to Deja Vu [36], in order to hide its cost. However, under batched inference, it is not always possible to fully mask the MLP router latency due to high GPU utilization. In practice, we observe that overlapping saves approximately 0.1 ms of router latency.
@@ Appendix D Sparse Kernels
1:Input: $A\in\mathbb{R}^{M\times K}$, $B\in\mathbb{R}^{K\times N}$, index vector $I$
@@
4:Gather indexed columns or rows of $B$ using $I$
@@ E.1 Pipeline Parallel Execution
Figure 5, 11 shows the decoding throughput for the OPT model family. For the OPT-6.7B model, despite around 90% sparsity in the MLP layer, the speedup is modest for small batch sizes due to insufficient workload to fully utilize the GPU’s parallel capabilities, highlighting the diminishing returns of sparsity in smaller models when the GPU is already underutilized. However, at larger batch sizes, speedups increase to 1.83×1.83$\times$1.83 × as attention computation becomes more memory-intensive. In contrast, the larger OPT-66B model shows more significant improvements, with speedups ranging from 1.66×1.$6$ at batch size 1 to $\times2$ at batch size 64, indicating better scaling with workload size.
Figure 6, 11 show the decoding throughput and speedup for the Llama model family. Since we sparsify only the attention layers, speedups are limited at smaller batch sizes due to the overhead of the router and the dominance of MLP layers in computational cost. As batch size increases, the computational burden shifts towards the attention layers, making sparsity more effective. For Llama 2 models, this results in speedups of up to 1.85×1.85$\times$1.85 ×.
@@ E.2 Decode Latency Analysis
Figure 13: OPT models decode, inter-token latency with fixed batch size of 16. Polar Sparsity reduces latency up to 2×2$\times$2 × compared to dense baseline and up to 1.52×1.52$\times$1.52 × compared to Deja Vu in this workload.
@@
Figure 14: LLaMA models decode, inter-token latency with fixed batch size of 16. Polar Sparsity reduces latency up to 1.77×1.$7$ compared to the dense baseline in this workload.
The figure 13 illustrates the decode latency of OPT 6.7b and OPT 66b models using dense, standard activation sparsity (Deja Vu), and our proposed Polar Sparsity methods across varying sequence lengths, with a fixed batch size of 16. Across both model scales, Polar Sparsity consistently achieves lower latency than both Dense and Deja Vu baselines, offering up to 2×2$\times$2 × speedup over Dense and up to 1.52×1.52$\times$1.52 × over Deja Vu in this workload. The figure 14 show the decode latency of LLaMA-2-7B and LLaMA-3.1-70B models across increasing sequence lengths using dense and Polar Sparsity, with batch size fixed to 16. Polar Sparsity consistently outperforms the dense baseline, achieving up to 1.77×1.$7$ speedup. As with throughput, even greater gains can be expected at larger batch sizes as the attention module dominates latency.

### v3

@@ C.1 Router Ablation Study
Figure 10 shows the decode latency of MLP and Attention blocks along with their respective routers, and includes the measured inference latency of sparse kernels across selected sparsity levels. The MLP router introduces approximately four times higher latency compared to the attention router. At higher activation levels, the combined latency of the MLP router and sparse MLP approaches, and in some cases exceeds, that of the dense MLP. To address this, we overlap the MLP router execution with the attention block, following a strategy similar to Deja Vu [Dejavu], in order to hide its cost. However, under batched inference, it is not always possible to fully mask the MLP router latency due to high GPU utilization. In practice, we observe that overlapping saves approximately 0.1 ms of router latency.
@@ Appendix D Sparse Kernels
1:Input: $A\in\mathbb{R}^{M\times K}$, $B\in\mathbb{R}^{K\times N}$, index vector II
@@
4:Gather indexed columns or rows of $B$ using II
@@
(a) Selective MLP Kernel
(b) Selective Attention Kernel
Figure 11: Selective MLP and Attention Kernel Performance across GPU Families. OPT 66b, batch size 64, sequence length 1920.
Figure 11 demonstrates that our selective MLP and attention kernels exhibit strong performance and deliver consistent speedups across NVIDIA GPU families (A5000, A100, and H100). As MLP and Attention density increases, latency scales near linearly, indicating that the kernels maintain high efficiency and effectively exploit sparsity without incurring significant overheads.
@@ E.1 Pipeline Parallel Execution
Figure 6, 12 shows the decoding throughput for the OPT model family. For the OPT-6.7B model, despite around 90% sparsity in the MLP layer, the speedup is modest for small batch sizes due to insufficient workload to fully utilize the GPU’s parallel capabilities, highlighting the diminishing returns of sparsity in smaller models when the GPU is already underutilized. However, at larger batch sizes, speedups increase to $1.83\times$ as attention computation becomes more memory-intensive. In contrast, the larger OPT-66B model shows more significant improvements, with speedups ranging from $1.66\times$ at batch size 1 to $2.2\times$ at batch size 64, indicating better scaling with workload size.
Figure 7, 12 show the decoding throughput and speedup for the Llama model family. Since we sparsify only the attention layers, speedups are limited at smaller batch sizes due to the overhead of the router and the dominance of MLP layers in computational cost. As batch size increases, the computational burden shifts towards the attention layers, making sparsity more effective. For Llama 2 models, this results in speedups of up to $1.85\times$.
@@ E.2 Decode Latency Analysis
Figure 14: OPT models decode, inter-token latency with fixed batch size of 16. Polar Sparsity reduces latency up to $2\times$ compared to dense baseline and up to $1.52\times$ compared to Deja Vu in this workload.
@@
Figure 15: LLaMA models decode, inter-token latency with fixed batch size of 16. Polar Sparsity reduces latency up to $1.77\times$ compared to the dense baseline in this workload.
The figure 14 illustrates the decode latency of OPT 6.7b and OPT 66b models using dense, standard activation sparsity (Deja Vu), and our proposed Polar Sparsity methods across varying sequence lengths, with a fixed batch size of 16. Across both model scales, Polar Sparsity consistently achieves lower latency than both Dense and Deja Vu baselines, offering up to $2\times$ speedup over Dense and up to $1.52\times$ over Deja Vu in this workload. The figure 15 show the decode latency of LLaMA-2-7B and LLaMA-3.1-70B models across increasing sequence lengths using dense and Polar Sparsity, with batch size fixed to 16. Polar Sparsity consistently outperforms the dense baseline, achieving up to $1.77\times$ speedup. As with throughput, even greater gains can be expected at larger batch sizes as the attention module dominates latency.